For each of the following situations give the degrees of freedom and an appropriate bound on the P-value (give the exact value if you have software available) for the χ2 statistic for testing the null hypothesis of no association between the row and column variables.

(a) A 2 by 2 table with χ2 = 1.08.

df =
P-value =


(b) A 4 by 4 table with χ2 = 18.83.

df =
P-value =


(c) A 2 by 8 table with χ2 = 22.43.

df =
P-value =


(d) A 5 by 3 table with χ2 = 12.62.

df =
P-value =
Answer #1

a) df = (2 - 1) * (2 - 1) = 1

p-value = 0.2987

b) df = (4 - 1) * (4 - 1) = 9

p-value = 0.0267

c) df = (2 - 1) * (8 - 1) = 7

p-value = 0.0021

d) df = (5 - 1) * (3 - 1) = 8

p-value = 0.1256